# Incremental Averaging Method to Improve Graph-Based Time-Difference-of-Arrival Estimation

### v2

@@ 2.2 Minimal Set Based on Common Reference Microphone
$\hat{m}_{\textrm{R}}\;=\;\operatorname*{argmax}_{m}\{\operatorname*{argmin}_{i\neq m}{\;R_{i,m}}\;\}\;.$ (6)
@@ 3.2 Incremental Method for Averaging CPSDs
Based on the vertices and edges of the MST, the proposed method builds up a consistent TDOA matrix with re-estimated TDOAs from scratch in $M-1$ steps, where $H$ denotes the step index. To improve the estimation of CPSDs corresponding to less reliable GCC-PHAT functions in later steps, we begin with CPSDs corresponding to the highest GCC-PHAT reliabilities. Therefore, before introducing the method, the edges are ordered based on reliability for which we define an $M-1$-dimensional vector $\mathbf{r}$ of reliabilities and a corresponding $(M-1)\times 2$-dimensional matrix $\mathbf{V}$ containing $M-1$ rows of MST edges $\left[i,j\right],\;i,j\in{\Theta}^{\textrm{MST}}$. The first entry of $\mathbf{r}$ contains the highest reliability value of the MST and the first row of $\mathbf{V}$ contains the pair of vertices connected by the associated edge. Subsequent entries of $\mathbf{r}$ can only correspond to an edge connected to one of the vertices from previous rows of $\mathbf{V}$, in order of descending reliability. For the example in Fig. 1, the reliability vector would be $\mathbf{r}_{\textrm{MST}}=[0.8,0.6,0.5,0.6]^{\textrm{T}}$ and the rows of $\mathbf{V}$ would be $[4,3]$, $[3,2]$, $[4,1]$, and $[5,1]$. The indices $i[H],\;\;j[H]$ refer to the microphones in the $H$-th row of $\mathbf{V}$. In each step $H$, the proposed method (referred to as MST+) re-estimates the TDOA between the microphones $i[H]$ and $j[H]$.
@@ 4.2 Practical Implementation
The algorithms were implemented with a sampling frequency $f_{s}=16$ kHz and a short-time Fourier transform (STFT) framework was used with a frame length of $1024$ samples (corresponding to 64 ms), 50% overlap between frames, and a square-root-Hann analysis window.
@@
To evaluate the usability of the estimated TDOAs from different methods for source position estimation, we used the spherical (SI) interpolation-based source position estimation method [1]. In each time-frame, the SI cost function (constrained to a distance of $5$ m from the centre of the room) was iteratively solved with gradient descent. For comparison, we also considered the SRP-PHAT method used in [11], i.e., the SRP-PHAT functional was evaluated on a 2D grid, first at a resolution of 10 cm, then at a resolution of 1 cm in the vicinities of the three grid points with the highest SRP-PHAT functional values, to find the global functional maximum.
@@ 4.3 Comparison of TDOA and Position Estimation Performance
Reverb. Error TDOA Estimation Method
@@
High $\overline{\sigma}$ [ms] 0.44 1.77 0.80 0.44 0.40 0.28
@@
For all considered TDOA estimation methods, Table 1 shows the TDOA and source position estimation errors for the three considered reverberation levels. As can be observed, the SRP-PHAT method results in a relatively poor source position estimation (and TDOA estimation) performance (e.g., $\overline{\sigma}=0.28$ ms, $\overline{\varepsilon}=29.3$ cm, and Acc$=63.0$ in medium reverberation). This is likely because the SRP-PHAT method has not been tested for such random source and microphone configurations, where it may not be ideal to weight the CPSD from each microphone pair equally within the SRP-PHAT functional. Considering the TDOA-based methods, randomly choosing the reference microphone (Ref-A method) also results in a relatively poor TDOA and source position estimation performance (e.g., $\overline{\sigma}=1.34$ ms, $\overline{\varepsilon}=70.9$ cm, and Acc$=54.3$ in medium reverberation). The results of the Ref-C method show that the TDOA and source position estimation performance can be easily improved by simply choosing the microphone closest to the centroid of the microphone array as the reference microphone (e.g., $\overline{\sigma}=0.54$ ms, $\overline{\varepsilon}=28.6$ cm, and Acc$=77.1$ in medium reverberation). The reliability-based Ref-R and MST methods outperform the SRP-PHAT, Ref-A and Ref-C methods, showing the importance of taking into account the reliability of the GCC-PHAT functions for determining the minimal set of microphone pairs. In low and medium reverberation conditions, the Ref-R and MST methods perform very similarly. However, in high reverberation the MST method performs slightly better than the Ref-R method ($\overline{\sigma}=0.40$ ms, $\overline{\varepsilon}=20.3$ cm, and Acc$=85.0$ compared to $\overline{\sigma}=0.44$ ms, $\overline{\varepsilon}=24.3$ cm, and Acc$=82.2$), most likely because in the MST method no reference microphone needs to be chosen. The results in Table 1 clearly show that the proposed MST+ method outperforms all other considered methods in every reverberation condition (e.g., $\overline{\sigma}=0.08$ ms, $\overline{\varepsilon}=5.0$ cm, and Acc$=95.8$ in medium reverberation). This suggests that it is beneficial to average over multiple CPSDs from multiple microphone pairs to accurately estimate TDOAs and improve source position estimation performance, rather than estimating each TDOA based on a single CPSD.

### v3

@@ 2.2 Minimal Set Based on Common Reference Microphone
$\hat{m}_{\textrm{R}}\;=\;\operatorname*{argmax}_{m}\{\operatorname*{argmin}_{i\neq m}{\;R_{i,m}}\;\}\;.\vskip-2.84526pt$ (6)
@@ 3.2 Incremental Method for Averaging CPSDs
Based on the vertices and edges of the MST, the proposed method builds up a consistent TDOA matrix with re-estimated TDOAs from scratch in $M-1$ steps, where $H$ denotes the step index. To improve the estimation of CPSDs corresponding to less reliable GCC-PHAT functions in later steps, we begin with CPSDs corresponding to the highest GCC-PHAT reliabilities. Therefore, before introducing the method, the edges are ordered based on reliability for which we define an $M-1$-dimensional vector $\mathbf{r}$ of reliabilities and a corresponding $(M-1)\times 2$-dimensional matrix $\mathbf{V}$ containing $M-1$ rows of MST edges $\left[i,j\right],\;i,j\in{\Theta}^{\textrm{MST}}$. The first entry of $\mathbf{r}$ contains the highest reliability value of the MST and the first row of $\mathbf{V}$ contains the pair of vertices connected by the associated edge. Subsequent entries of $\mathbf{r}$ can only correspond to an edge connected to one of the vertices from previous rows of $\mathbf{V}$, in order of descending reliability. For the example in Fig. 1, the reliability vector would be $\mathbf{r}_{\textrm{MST}}=[0.8,0.6,0.5,0.6]^{\textrm{T}}$ and the rows of $\mathbf{V}$ would be [4,3][4,3][ 4 , 3 ], [3,2][3,2][ 3 , 2 ], [4,1][4,1][ 4 , 1 ], and [5,1][5,1][ 5 , 1 ]. The indices $i[H],\;\;j[H]$ refer to the microphones in the $H$-th row of $\mathbf{V}$. In each step $H$, the proposed method (referred to as MST+) re-estimates the TDOA between the microphones $i[H]$ and $j[H]$.
@@ 4.2 Practical Implementation
The algorithms were implemented with a sampling frequency $f_{s}=16$ kHz and a short-time Fourier transform (STFT) framework was used with a frame length of 102410241024 samples (corresponding to 64 ms), 50% overlap between frames, and a square-root-Hann analysis window.
@@
To evaluate the usability of the estimated TDOAs from different methods for source position estimation, we used the spherical (SI) interpolation-based source position estimation method [1]. In each time-frame, the SI cost function (constrained to a distance of 555 m from the centre of the room) was iteratively solved with gradient descent. For comparison, we also considered the SRP-PHAT method used in [11], i.e., the SRP-PHAT functional was evaluated on a 2D grid, first at a resolution of 10 cm, then at a resolution of 1 cm in the vicinities of the three grid points with the highest SRP-PHAT functional values, to find the global functional maximum.
@@ 4.3 Comparison of TDOA and Position Estimation Performance
Reverb.
Error TDOA Estimation Method
@@
High
$\overline{\sigma}$ [ms] 0.44 1.77 0.80 0.44 0.40 0.28
@@
For all considered TDOA estimation methods, Table 1 shows the TDOA and source position estimation errors for the three considered reverberation levels. As can be observed, the SRP-PHAT method results in a relatively poor source position estimation (and TDOA estimation) performance (e.g., $\overline{\sigma}=0.28$ ms, $\overline{\varepsilon}=29.3$ cm, and Acc=63.0=63.0= 63.0 % in medium reverberation). This is likely because the SRP-PHAT method has not been tested for such random source and microphone configurations, where it may not be ideal to weight the CPSD from each microphone pair equally within the SRP-PHAT functional. Considering the TDOA-based methods, randomly choosing the reference microphone (Ref-A method) also results in a relatively poor TDOA and source position estimation performance (e.g., $\overline{\sigma}=1.34$ ms, $\overline{\varepsilon}=70.9$ cm, and Acc=54.3=54.3= 54.3 % in medium reverberation). The results of the Ref-C method show that the TDOA and source position estimation performance can be easily improved by simply choosing the microphone closest to the centroid of the microphone array as the reference microphone (e.g., $\overline{\sigma}=0.54$ ms, $\overline{\varepsilon}=28.6$ cm, and Acc=77.1=77.1= 77.1 % in medium reverberation). The reliability-based Ref-R and MST methods outperform the SRP-PHAT, Ref-A and Ref-C methods, showing the importance of taking into account the reliability of the GCC-PHAT functions for determining the minimal set of microphone pairs. In low and medium reverberation conditions, the Ref-R and MST methods perform very similarly. However, in high reverberation the MST method performs slightly better than the Ref-R method ($\overline{\sigma}=0.40$ ms, $\overline{\varepsilon}=20.3$ cm, and Acc=85.0=85.0= 85.0 % compared to $\overline{\sigma}=0.44$ ms, $\overline{\varepsilon}=24.3$ cm, and Acc=82.2=82.2= 82.2 %), most likely because in the MST method no reference microphone needs to be chosen. The results in Table 1 clearly show that the proposed MST+ method outperforms all other considered methods in every reverberation condition (e.g., $\overline{\sigma}=0.08$ ms, $\overline{\varepsilon}=5.0$ cm, and Acc=95.8=95.8= 95.8 % in medium reverberation). This suggests that it is beneficial to average over multiple CPSDs from multiple microphone pairs to accurately estimate TDOAs and improve source position estimation performance, rather than estimating each TDOA based on a single CPSD.